1. For the following equation, find the center, vertices, foci, transverse axis, and asymptotes, and sketch the graph: 2. Consider the set of parametric equations (a) Graph in the following window: TMIN--3.74, TMAX- 3.74, TSTEP = 0.02, XMIN =-10, XMAX = 10, YMIN =-7, YMAX = 7, Sketch the graph. (b) At, find (x, y) and dy/dx. Write the equations of the lines tangent to and normal to the graph at (c) Find the length of the curve from to
2 Consider the set of parametric equations (a) Graph in the following window: TMIN -3.74, TMAX-3.74, TSTEP = 0.02, XMIN =-10, XMAX = 10, Y MIN =-7, YMAX = 7, Sketch the graph. (b) At, find (x, y) and dy/dx. Write the equations of the lines tangent to and normal to the graph at (c) Find the length of the curve from to (d) Find the area of the surface resulting from revolving the length of the curve in (c) about the x-axis. (e) Find the area of the surface resulting from revolving the length of the curve in (c) about the y-axis.
1. For the following equation, find the center, vertices, foci, transverse axis, and asymptotes, and sketch the graph: 4x2-y2 + 4x + 4y + 1 = 0. 2. Consider the set of parametric equations = t + cos(nt) (a) Graph in the following window: TMIN--3.74, TMAX 3.74, TSTEP YMIN =-7, YMAX 7. Sketch the graph. 0.02, XMIN -10, XMAX 10,
(b) At t = 1.5, find (x, y) and dy/dx. Write the equations of the lines tangent to and normal to the graph at (c) Find the length of the curve from t 3.1tot 3.7. (d) Find the area of the surface resulting from revolving the length of the curve in (c) about the x-axis. (e) Find the area of the surface resulting from revolving the length of the curve in (c) about the y-axis.
Answer #1

(1)Given 4x2-y2+4x+4y+1-0 Complete the square first 4 +x+(3) Ho,2-4y + 4) + 4 = 0 1+1-(y-2)2 + 4 = 0 (y-2)-4x+4-4 4 4 (y-2) 4